# ACT: Agentic Classification Tree

### v2

@@ 3.2 Proposed Approach: Agentic Classification Tree (ACT)
  1. 1.
    Quantitative Impurity Evaluation. As in classical decision tree algorithms such as CART and C4.5, the quality of a split induced by the current prompt $p^{(k)}$ is evaluated using standard impurity criteria (e.g., weighted Gini impurity (Breiman et al., 1984) or information gain ratio (Salzberg, 1994)). We denote this score by $\delta(p^{(k)})$, which serves as an objective function quantifying how well the semantic split separates the class labels.
  2. 2.
    Semantic Purity Analysis via LLM. While the quantitative evaluation (e.g., Gini impurity) measures class-label heterogeneity within each partition, it does not reveal the semantic factors underlying the impurity. To make these sources explicit, we analyze each child node independently by contrasting its class-conditional subsets (correct vs. misclassified examples). For this purpose, we align the predicted answer (yes/no) with the ground-truth label ($y_{i}\in\{0,1\}$) when defining correct and incorrect routing. Although this introduces an artificial correspondence between branch outcomes and class labels, it proved to be more effective in practice: the LLM more readily exploits feedback framed in terms of correct versus incorrect predictions than feedback based on contrasting the two branches directly (e.g., retaining the majority class and pruning the minority one). This alignment does not prevent the emergence of negative queries (e.g., “Is feature $A$ absent?”), which the LLM can in principle generate and refine, though such formulations are typically harder to elicit consistently.
    Formally, each subset $g\in\{\mathcal{D}_{L}^{x},\mathcal{D}_{R}^{x}\}$ is associated with the model’s predicted answer $\hat{y}_{g}$ to the current prompt. For convenience, we denote by $\hat{y}_{g}\in\{0,1\}$ the numeric encoding of this answer ($\hat{y}_{g}=1$ for “yes”, $\hat{y}_{g}=0$ for “no”). We then further partition each subset into two groups:
    $X_{\text{correct}}^{x}=\{(x_{i},y_{i})\in g\mid y_{i}=\hat{y}_{g}\},\quad X_{\text{error}}^{x}=\{(x_{i},y_{i})\in g\mid y_{i}\neq\hat{y}_{g}\}.$
@@
    Next, we prompt the LLM to analyze these two groups and identify key semantic characteristics that distinguish them—i.e., features whose presence or absence could explain why some examples are misclassified.
    Based on this contrast, the LLM—denoted $f_{\text{purity}}$—returns concise, actionable feedback $s_{g}$ that is then used to refine the node’s splitting question in the next optimization step to reduce impurity.
    LLM Task: Semantic Purity Feedback for "yes" group ($D_{L}$) Below are two groups of samples for which a model answered either "yes" or "no" in response to a natural language prompt to predict their class label. Provide feedback about key characteristics that are present or absent in the group where the model’s predicted label matched the true label, and in the group where the prediction was incorrect. For the following examples, the model answered "yes": Well-classified examples (true label = "yes"):
@@
    By performing this semantic analysis independently on subsets $D_{L}$ and $D_{R}$, we obtain targeted feedback $s_{D_{L}}$ and $s_{D_{R}}$ that are then used to guide subsequent question-refinement iterations.
Best split search
At each decision node, the objective is to refine the natural-language prompt $p^{(k)}$ so that the resulting partition minimizes the weighted Gini impurity $\delta(p^{(k)})$. As described previously, this quantitative criterion is complemented by semantic feedback $(s_{D_{L}},s_{D_{R}})$, obtained from LLM-based analyses of misclassified versus correctly classified examples.
To jointly evaluate statistical impurity and semantic error patterns, we define a semantic–impurity objective:
@@
where $f_{\mathrm{loss}}$ is instantiated as a large language model (LLM). Given the current question, impurity statistics, and semantic feedback, the LLM returns a natural-language diagnosis of its limitations, highlighting semantic factors that contribute to impurity.
We realize the refinement process using the differentiable prompting framework TextGrad (Yuksekgonul et al., 2024). Each refinement iteration proceeds in two stages:
  1. (i)
    Guidance stage: TextGrad.feedback takes as input the prompt $p^{(k)}$ together with its evaluation $\mathcal{L}(p^{(k)})$, and generates a natural-language editing instruction $\nabla_{p}\mathcal{L}(p^{(k)})$ that specifies how the prompt should be revised.
  2. (ii)
@@
The optimization loop is repeated for up to $K$ steps or until convergence. At each iteration, the current prompt $p^{(k)}$ is used to partition the data, evaluate impurity, and analyze semantic error patterns via the LLM, as described above. These signals jointly inform the loss $\mathcal{L}(p^{(k)})$, which is then used by TextGrad to generate a refined prompt $p^{(k+1)}$. Among all prompts generated during the optimization process, $p^{(0)},p^{(1)},\dots,p^{(K)}$, we select the one with the lowest impurity on the training set, i.e., $p^{\ast}=\arg\min_{k}\,\delta(p^{(k)})$.
Tree construction
ACT builds the decision tree recursively in a top-down manner. At each node, the best prompt is optimized via the procedure described above. If the resulting subset is pure or meets a predefined stopping criterion (e.g., minimum node size — the number of training examples in the node —, Gini impurity below a threshold, or maximum depth), a leaf node is created. As in classical decision trees, the leaf is assigned the majority class label among the examples it contains. The complete ACT procedure is summarized in Algorithm 1.
@@
$T\leftarrow$ GrowPromptTree(D)(D) ;
return $T$ ;
Function GrowPromptTree($D$):
   Initialize prompt $p^{(0)}$ ;
 // default initialization
@@
    return Leaf node with majority class label;
@@
    $D_{L}\leftarrow\{(x_{i},y_{i})\in D\mid f_{\text{split}}(p^{(k)},x_{i})=\texttt{yes}\}$;
    $D_{R}\leftarrow D\setminus D_{L}$;
    $\delta(p^{(k)})\leftarrow\frac{|D_{L}|}{|D|}\cdot\mathrm{Gini}(D_{L})+\frac{|D_{R}|}{|D|}\cdot\mathrm{Gini}(D_{R})$ ;
    foreach $(g,\hat{y}_{g})\in\{(D_{L},\texttt{'yes'}),(D_{R},\texttt{'no'})\}$ do
       $X_{\text{correct}}\leftarrow\{(x_{i},y_{i})\in g\mid y_{i}=\hat{y}_{g}\}$ ;
       $X_{\text{error}}\leftarrow\{(x_{i},y_{i})\in g\mid y_{i}\neq\hat{y}_{g}\}$ ;
       $s_{g}\leftarrow f_{\text{purity}}(p^{(k)},X_{\text{correct}},X_{\text{error}})$ ;
       // Intra-group semantic analysis
    $\mathcal{L}(p^{(k)})\leftarrow f_{\text{loss}}(p^{(k)},s_{D_{L}},s_{D_{R}},\delta(p^{(k)}))$ ;
    // LLM-evaluated semantic loss
    $\nabla_{p}\mathcal{L}(p^{(k)})\leftarrow\texttt{TextGrad.feedback}(p^{(k)},\mathcal{L}(p^{(k)}))$ ;
    $p^{(k+1)}\leftarrow\texttt{TextGrad.step}(p^{(k)},\nabla_{p}\mathcal{L}(p^{(k)}))$ ;
 $p^{\ast}\leftarrow\arg\min_{p^{(k)}}\delta(p^{(k)})$ ;
 // Prompt with lowest impurity
   Split $D$ using $p^{\ast}$ into $D_{L}$ and $D_{R}$ ;
   node $\leftarrow$ Create prompt node with final question $p^{\ast}$ ;
@@
 return node ;
Function Gini($D$):
 return $1-\sum_{k}p_{k}^{2}$, where $p_{k}$ is the class proportion in $D$;
Algorithm 1 ACT: Agentic Classification Tree
Table 2: Comparison of classification methods across three text datasets (DIAGNO, SPAM, JAILBREAK) and four LLMs. Results show training and test accuracy (%) for baseline methods (CoT, DSPy, TextGrad), traditional CART with TF-IDF features, and the proposed ACT with varying hyperparameters ($d$: tree depth, $k$: optimization steps per node).
Dataset Method Gemma3 4b GPT-4.1 Nano GPT-4.1 Mini Qwen3 4b Avg Test
Train Test Train Test Train Test Train Test (%)
DIAGNO CoT (0-shot) 61.0 61.5 63.5 63.3 64.3 61.2 63.5 63.2 62.3
DSPy (8 demos) 66.2 64.0 71.5 68.7 68.3 64.5 61.2 58.8 64.0
TextGrad 65.7 63.3 67.5 64.5 69.2 65.8 65.0 64.0 64.4
TF-IDF + CART (d=3d{=}3) 78.8
ACT (d=3d{=}3, k=10k{=}10) 71.3 65.3 70.7 66.9 78.5 77.3 68.5 64.8 68.6
ACT (d=3d{=}3, k=20k{=}20) 72.5 67.0 71.8 67.2 78.3 76.2 69.2 65.5 69.0
ACT (d=4d{=}4, k=10k{=}10) 72.5 66.3 74.3 70.5 79.7 74.3 70.8 66.7 69.5
ACT (d=4d{=}4, k=20k{=}20) 73.7 68.3 74.8 70.3 85.0 82.3 75.5 70.3 72.8
SPAM CoT (0-shot) 75.0 73.3 96.3 95.5 95.5 95.7 92.5 93.2 89.4
DSPy (8 demos) 95.2 95.0 98.3 98.3 98.8 98.2 96.7 96.5 97.0
TextGrad 93.3 92.6 97.8 97.3 98.0 96.7 96.2 96.0 95.7
TF-IDF + CART (d=3d{=}3) 89.3
ACT (d=2d{=}2, k=5k{=}5) 97.0 95.8 98.7 98.6 98.5 98.2 97.1 96.6 97.3
ACT (d=2d{=}2, k=10k{=}10) 97.7 96.5 99.7 99.2 100.0 99.2 98.6 98.5 98.4
ACT (d=3d{=}3, k=5k{=}5) 98.3 98.2 98.3 97.0 100.0 99.4 98.8 98.5 98.3
ACT (d=3d{=}3, k=10k{=}10) 99.3 98.5 99.2 98.8 100.0 99.1 99.0 98.8 98.8
JAILBREAK CoT (0-shot) 81.6 77.9 84.1 85.5 88.1 90.4 78.0 80.1 83.5
DSPy (8 demos) 85.8 79.8 88.0 88.0 94.8 94.8 85.3 83.5 86.5
TextGrad 92.9 91.3 92.4 88.2 95.6 94.7 82.3 81.5 88.9
TF-IDF + CART (d=3d{=}3) 91.3
ACT (d=3d{=}3, k=10k{=}10) 85.3 82.3 85.5 85.5 96.9 95.4 84.8 82.7 86.5
ACT (d=3d{=}3, k=20k{=}20) 85.2 84.3 92.3 92.0 97.7 96.2 84.2 84.7 89.3
ACT (d=4d{=}4, k=10k{=}10) 91.4 90.0 92.7 87.0 98.1 97.2 85.8 85.1 89.8
ACT (d=4d{=}4, k=20k{=}20) 92.1 90.6 93.2 88.2 99.2 98.9 87.7 88.0 91.4
@@ 4 Experiments
We empirically evaluate the proposed Agentic Classification Tree (ACT) against state-of-the-art baseline methods on multiple binary text classification datasets.
Datasets. We evaluate ACT on three binary classification tasks: medical diagnosis (DIAGNO), spam detection (SPAM), and jailbreak prompt classification (JAILBREAK). For DIAGNO and SPAM, we construct balanced splits to ensure controlled evaluation. JAILBREAK is used in its original form, which is moderately imbalanced (details in Appendix B.1).

### v3

@@ 3.2 Proposed Approach: Agentic Classification Tree
(1) Quantitative Impurity Evaluation. As in classical decision tree algorithms, the quality of a split induced by the current prompt $p^{(k)}$ is evaluated using standard impurity criteria $\delta(p^{(k)})$, e.g. weighted Gini impurity, quantifying how well the semantic split separates the class labels.
(2) Semantic Purity Analysis via LLM. While the quantitative evaluation (e.g., Gini impurity) measures class-label heterogeneity within each partition, it does not reveal the semantic factors underlying the impurity. To make these sources explicit, we analyze each child node independently by contrasting its class-conditional subsets (correct vs. misclassified examples). For this purpose, we align the predicted answer (yes/no) with the ground-truth label ($y_{i}\in\{0,1\}$) when defining correct and incorrect routing. Although this introduces an artificial correspondence between branch outcomes and class labels, it proved to be more effective in practice: the LLM more readily exploits feedback framed in terms of correct versus incorrect predictions than feedback based on contrasting the two branches directly (e.g., retaining the majority class and pruning the minority one). This alignment does not prevent the emergence of negative queries (e.g., “Is feature $A$ absent?”), which the LLM can in principle generate and refine, though such formulations are typically harder to elicit consistently. Formally, each subset $g\in\{\mathcal{D}_{L}^{n},\mathcal{D}_{R}^{n}\}$ is associated with the model’s predicted answer $\hat{y}_{g}$ to the current prompt. For convenience, we denote by $\hat{y}_{g}\in\{0,1\}$ the numeric encoding of this answer ($\hat{y}_{g}=1$ for “yes”, $\hat{y}_{g}=0$ for “no”). We then further partition each subset into two groups:
$X_{\text{correct}}^{n}$ $=\{(x_{i},y_{i})\in g\mid y_{i}=\hat{y}_{g}\},$
$X_{\text{error}}^{n}$ $=\{(x_{i},y_{i})\in g\mid y_{i}\neq\hat{y}_{g}\}.$
@@
Next, we prompt the LLM to analyze these two groups and identify key semantic characteristics that distinguish them—i.e., features whose presence or absence could explain the misclassification of some examples. Based on this contrast, the LLM (denoted $f_{\text{purity}}$) returns concise, actionable feedback $s_{g}$ that is then used to refine the node’s splitting question in the next optimization step to reduce impurity. By performing this semantic analysis independently on subsets $D_{L}$ and $D_{R}$, we obtain targeted feedback $s_{D_{L}}$ and $s_{D_{R}}$ that are then used to guide subsequent question-refinement iterations.
LLM Task: Semantic Purity Feedback for "yes" group ($D_{L}$) Below are two groups of samples for which a model answered either "yes" or "no" in response to a natural language prompt to predict their class label. Provide feedback on key characteristics that are present or absent in the group where the model’s predicted label matched the true label, and in the group where the prediction was incorrect. For the following examples, the model answered "yes": Well-classified examples (true label = "yes"):
@@
Best split search.
At each node, the objective is to refine the prompt $p^{(k)}$ so that the resulting partition minimizes the weighted Gini impurity $\delta(p^{(k)})$. As described previously, this quantitative criterion is complemented by semantic feedback $(s_{D_{L}},s_{D_{R}})$, obtained from LLM-based analyses of misclassified versus correctly classified examples. To jointly evaluate statistical impurity and semantic error patterns, we define a semantic–impurity objective:
@@
where $f_{\mathrm{loss}}$ is instantiated as an LLM. Given the current question, impurity statistics and semantic feedback, the LLM returns a natural-language diagnosis of its limitations, highlighting semantic factors that contribute to impurity.
We realize the refinement process using the differentiable prompting framework TextGrad (yuksekgonul2024textgrad). Each refinement iteration proceeds in two stages:
  1. 1.
    Guidance stage: TextGrad.feedback takes the prompt $p^{(k)}$ and its evaluation $\mathcal{L}(p^{(k)})$ as input, and generates a natural-language editing instruction $\nabla_{p}\mathcal{L}(p^{(k)})$ specifying how to revise the prompt.
  2. 2.
@@
The optimization loop runs for up to $K$ steps or until convergence. At each iteration, the current prompt $p^{(k)}$ is used to partition the data, evaluate impurity and analyze semantic error patterns via the LLM, as described above. These signals jointly inform the loss $\mathcal{L}(p^{(k)})$, which is then used by TextGrad to generate a refined prompt $p^{(k+1)}$. Among all candidate prompts $p^{(0)},\dots,p^{(K)}$, we select the one with the lowest training impurity $p^{\ast}=\arg\min_{k}\,\delta(p^{(k)})$.
Tree construction.
As summarized in Algorithm 1, ACT builds the decision tree recursively in a top-down manner. At each node, the best prompt is optimized via the procedure described above. If the resulting subset is pure or meets a predefined stopping criterion (e.g., minimum node size, Gini impurity below a threshold, or maximum depth), a leaf node is created. As in classical decision trees, the leaf is assigned the majority class label among the examples it contains.
Tree interpretability.
Decision trees, which rely on transparent structures and parameters to make decisions, are widely considered to be interpretable (molnar2020interpretable). Owing to its shallow decision-tree structure and natural-language node questions, ACT exhibits the same desirable properties. In practice, interpretability depends on both the tree depth (i.e., the number of questions along a decision path) and the clarity of the questions. Accordingly, we restrict the tree depth to 3–4 to keep decision paths concise and explicitly instruct TextGrad during training to generate simple questions and limiting the number of logical clauses (and/or) to at most $L=2$, as detailed in Appendix B.2.
Algorithm 1 ACT: Agentic Classification Tree
@@
function GrowPromptTree($D$)
  Initialize prompt $p^{(0)}$ {default initialization}
@@
    return Leaf node with majority class label
  end if
@@
    $D_{L}\leftarrow\{(x_{i},y_{i})\in D\mid f_{\text{split}}(p^{(k)},x_{i})=\texttt{yes}\}$
    $D_{R}\leftarrow D\setminus D_{L}$
    $\delta(p^{(k)})\leftarrow\frac{|D_{L}|}{|D|}\,\mathrm{Gini}(D_{L})\;+\;\frac{|D_{R}|}{|D|}\,\mathrm{Gini}(D_{R})$
    for all $(g,\hat{y}_{g})\in\{(D_{L},\texttt{yes}),(D_{R},\texttt{no})\}$ do
     $X_{\text{correct}}\leftarrow\{(x_{i},y_{i})\in g\mid y_{i}=\hat{y}_{g}\}$
     $X_{\text{error}}\leftarrow\{(x_{i},y_{i})\in g\mid y_{i}\neq\hat{y}_{g}\}$
     $s_{g}\leftarrow f_{\text{purity}}(p^{(k)},X_{\text{correct}},X_{\text{error}})$
    end for
    $\mathcal{L}(p^{(k)})\leftarrow f_{\text{loss}}(p^{(k)},s_{D_{L}},s_{D_{R}},\delta(p^{(k)}))$
@@
  end for
  $p^{\ast}\leftarrow\arg\min_{p^{(k)}}\delta(p^{(k)})$ {lowest impurity}
  Split $D$ using $p^{\ast}$ into $D_{L}$ and $D_{R}$
  node $\leftarrow$ Create prompt node with final question $p^{\ast}$
  node.left $\leftarrow$ GrowPromptTree($D_{L}$)
  node.right $\leftarrow$ GrowPromptTree($D_{R}$)
  return node
end function
function Gini($D$)
  return $1-\sum_{k}p_{k}^{2}$ $p_{k}$ = class proportion in DD}
end function
@@ 4 Experiments
Table 2: Comparison of classification methods across five datasets and four LLMs. Results show accuracy and F1-score on the test set (in %) for six baselines, as well as the proposed ACT with varying hyperparameters ($d$: tree depth, $k$: optimization steps per node).
Method Gemma3 4b GPT-4.1 Nano GPT-4.1 Mini Qwen3 4b Avg
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
DIAGNO
CoT (0-shot) 61.5 69.5 63.3 54.7 61.2 51.4 63.2 41.7 62.3 54.3
DSPy (BFSR, 8 demos) 64.0 68.2 68.7 61.2 64.5 59.4 63.3 54.9 65.1 60.9
TextGrad 63.3 56.7 64.5 53.0 65.8 55.5 64.0 60.0 64.4 56.3
TF-IDF + CART (d=3d{=}3) -- -- -- -- -- -- -- -- $78.8$ $78.9$
Con. tag. 65.5 63.9 59.3 58.1 75.3 78.6 58.8 64.7 64.7 66.3
Rule Fit -- -- -- -- -- -- -- -- $80.2$ $79.3$
ACT (d=3d{=}3, k=10k{=}10) 65.3 66.5 66.8 67.0 77.3 75.2 64.8 61.3 68.6 67.5
ACT (d=4d{=}4, k=20k{=}20) 68.3 70.8 70.3 70.8 82.3 81.8 70.3 73.5 72.8 74.2
SPAM
CoT (0-shot) 73.3 78.6 95.5 95.3 95.7 95.5 93.2 93.6 89.4 90.8
DSPy (BFSR, 8 demos) 95.0 95.2 98.3 98.3 98.2 98.2 97.2 97.2 97.2 97.2
TextGrad 95.0 94.7 97.3 97.3 96.7 96.6 96.0 96.0 96.3 96.2
TF-IDF + CART (d=5d{=}5) -- -- -- -- -- -- -- -- $92.7$ $93.1$
Con. tag. 83.2 85.5 64.2 45.8 82.3 85.0 62.3 39.6 73.0 64.0
Rule Fit -- -- -- -- -- -- -- -- $97.0$ $96.9$
ACT (d=2d{=}2, k=5k{=}5) 95.8 95.8 98.7 98.6 98.2 98.2 96.7 96.6 97.4 97.3
ACT (d=3d{=}3, k=10k{=}10) 98.5 98.5 99.2 99.2 99.5 99.5 98.8 98.8 99.0 99.0
JAILBREAK
CoT (0-shot) 77.9 82.2 85.5 83.5 93.6 93.3 81.1 77.7 84.5 84.2
DSPy (BFSR, 8 demos) 89.6 89.3 91.6 91.6 94.8 94.8 85.5 83.9 90.4 89.9
TextGrad 91.6 91.3 90.4 90.3 94.8 95.1 88.8 90.0 91.4 91.7
TF-IDF + CART (d=4d{=}4) -- -- -- -- -- -- -- -- $92.8$ $92.4$
Con. tag. 82.7 82.2 58.6 31.8 82.7 79.8 76.3 70.1 75.1 66.0
Rule Fit -- -- -- -- -- -- -- -- $96.8$ $96.8$
ACT (d=3d{=}3, k=10k{=}10) 82.3 82.7 85.5 85.1 95.2 95.5 82.7 82.6 86.4 86.5
ACT (d=4d{=}4, k=20k{=}20) 90.8 90.0 92.0 91.5 98.8 98.8 92.0 92.7 93.4 93.3
BANKCHURN
CoT (0-shot) 47.8 48.1 48.5 25.9 52.8 60.5 50.2 37.3 49.8 43.0
DSPy (BFSR, 8 demos) 52.8 55.6 52.3 42.8 57.2 58.5 54.0 52.2 54.1 52.3
TextGrad 52.0 60.0 53.3 60.3 55.3 61.0 52.8 61.4 53.4 60.7
TF-IDF + CART (d=5d{=}5) -- -- -- -- -- -- -- -- $63.2$ $66.8$
Con. tag. 57.2 60.2 52.0 59.8 59.7 60.8 54.3 61.6 55.8 60.6
Rule Fit -- -- -- -- -- -- -- -- $63.0$ $59.8$
ACT (d=3d{=}3, k=10k{=}10) 56.3 42.3 58.8 67.4 60.5 64.6 57.5 62.7 58.3 59.3
ACT (d=4d{=}4, k=20k{=}20) 65.2 70.7 65.0 69.1 68.7 68.6 62.0 60.0 65.2 67.1
IMDB
CoT (0-shot) 92.3 91.9 94.2 94.1 95.7 95.6 94.2 94.0 94.1 93.9
DSPy (BFSR, 8 demos) 93.7 93.6 95.5 95.5 96.8 96.8 94.8 94.8 95.2 95.2
TextGrad 93.2 92.9 95.2 95.1 95.8 95.8 94.7 94.5 94.7 94.6
TF-IDF + CART (d=3d{=}3) -- -- -- -- -- -- -- -- $66.8$ $72.8$
Con. tag. 74.8 76.9 56.2 31.7 50.0 66.7 51.3 63.8 58.1 59.8
Rule Fit -- -- -- -- -- -- -- -- $75.2$ $73.7$
ACT (d=3d{=}3, k=10k{=}10) 93.2 93.2 92.8 92.8 95.7 95.7 93.8 93.9 93.9 93.9
ACT (d=3d{=}3, k=20k{=}20) 93.7 93.6 94.7 94.7 96.3 96.2 95.5 95.5 95.1 95.0
@@
We empirically evaluate ACT against state-of-the-art baseline methods on multiple binary text classification datasets.
4.1 Setup: Datasets, Models & Baselines
Datasets.
We evaluate ACT on five binary text-classification tasks: The DIAGNO dataset contains short clinical descriptions labeled as tuberculosis (TB) or not TB, and we use balanced train–test splits for controlled evaluation. SPAM consists of email messages annotated as spam or ham, also evaluated with balanced splits. JAILBREAK contains user prompts labeled according to whether they attempt to circumvent safety policies, and is used in its original moderately imbalanced form. BANKCHURN is a customer-churn dataset in which tabular attributes are serialized into short textual profiles following the protocol used by (arzaghi2025intrinsic). This task emphasizes dataset-specific attribute relationships rather than broad world knowledge. IMDB is a standard sentiment dataset of movie reviews; we use a balanced subset for consistency.